# No Nowhere Constant Continuous Function Maps All Non-Normal Numbers to Normal Numbers

### v3

@@ 2 Proof of Main Results
We note that the size of $\Omega_{M}^{\varepsilon}$ can be written in terms of a binomial expansion
$|\Omega_{M}^{\varepsilon}|=\sum_{K\geq(1-\varepsilon)M}^{M}\binom{M}{K}(b-1)^{M-K}\leq\varepsilon M\binom{M}{\lceil(1-\varepsilon)M\rceil}(1-b)^{\varepsilon N},$ (A.3)
@@
for any integer $K$ and any $\varepsilon>0.$ Moreover, we note that
@@ Acknowledgments
The author would like to thank Verónica Becher and Christoph Aistleitner for their support and for providing valuable references. Special thanks are due to the anonymous referee for numerous helpful comments that substantially improved the paper; in particular, the referee inspired Theorem 1.5 and provided the proof of its second statement.

### v4

@@ Appendix A Proof of $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$
The size of $\Omega_{M}^{\varepsilon}$ can be written in terms of a binomial expansion
$|\Omega_{M}^{\varepsilon}|=\sum_{K\geq(1-\varepsilon)M}^{M}\binom{M}{K}(b-1)^{M-K}\leq\varepsilon M\binom{M}{\lceil(1-\varepsilon)M\rceil}(1-b)^{\varepsilon M},$ (A.4)
@@
for any integer $K$ and any $\varepsilon>0.$ Moreover,
@@
Appendix B Simultaneous non-normality for different bases
In this appendix, we discuss the non-normality for various bases.
Corollary B.1.
$\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset$ (B.1)
Proof.
@@ Acknowledgments
The author would like to thank Verónica Becher and Christoph Aistleitner for their support and for providing valuable references. Special thanks are due to the anonymous referee for numerous helpful comments on the structure and the content of this work that substantially improved the paper. In particular, the referee inspired Theorem 1.3 and provided proof ideas for its first statement. More precisely, the proofs of Proposition 2.3 and Corollary 2.4 follow his suggestions.